# with Instruction Back-Translation

### v2

@@ Abstract
In this paper, we introduce KunThe dataset is named Kun as Chinese pronunciation of Humpback Li et al. (2023c)., a novel approach for creating high-quality instruction-tuning datasets for large language models (LLMs) without relying on manual annotations. Adapting a self-training algorithm based on instruction back-translation and answer polishment, Kun leverages unlabelled data from diverse sources such as Wudao, Wanjuan, and SkyPile to generate a substantial dataset of over a million Chinese instructional data points. This approach presents a novel departure from traditional methods by using a self-curation process to refine and select the most effective instruction-output pairs. Our experiments with the 6B-parameter Yi model across various benchmarks demonstrate Kun’s robustness and scalability. Our method’s core contributions lie in its algorithmic advancement, which enhances data retention and clarity, and its innovative data generation approach that substantially reduces the reliance on costly and time-consuming manual annotations. This methodology presents a scalable and efficient solution for improving the instruction-following capabilities of LLMs, with significant implications for their application across diverse fields. The code and dataset can be found at https://github.com/Zheng0428/COIG-Kun
Kun: Answer Polishment for Chinese Self-Alignment
with Instruction Back-Translation
[…]
footnotetext: *  These authors contributed equally.footnotetext:   Corresponding Authors.
@@ 1 Introduction
AP is central to Kun’s strategy. It addresses a critical challenge in the Humpback (Li et al., 2023c) method, where raw data, once labeled, are directly used in instruction datasets. The unscreened raw data often mismatches between instructions and responses, as raw data may not inherently align with the instructional context. AP refines this raw data, ensuring a tighter correlation between the instructions and responses through a back-translation process. This leads to a dataset where each instruction-output pair is more coherent and contextually relevant. Unlike methods dependent on LLMs (Peng et al., 2023; Taori et al., 2023a; Zheng et al., 2023), Kun offers an independent and scalable approach to instruction-based training.
We opt for the 6B-parameter Yi model due to its open-source nature and dependable performancehttps://github.com/01-ai/Yi. Its efficacy is tested and proven across diverse dataset sizes, including widely recognized benchmarks like C-EVAL (Huang et al., 2023) and CMMLU (Li et al., 2023b). To evaluate the performance of the model, we design acomprehensive human evaluation which contains 500 prompts from ShareGPT-zh, covering various tasks. Responses generated by our model are compared with those from other models, showcasing the superiority of our Kun-52k variant. Further details can be found in 4.3.1. Additionally, we evaluate the quality of our dataset, which includes 1,000 instruction-output pairs each from sources like Wudao (Xue et al., 2022), Wanjuan (He et al., 2023), and SkyPile (Wei et al., 2023). This evaluation, focusing on clarity, feasibility, practicality, and alignment, ensures the high quality of our dataset. The key contributions of our work are:
@@ 2 Related Work
Instruction Tuning Instruction tuning is widely recognized as a key technique for activating LLMs to adhere to human conversational norms.  (Mishra et al., 2022; Wang et al., 2022b, 2023b). Instruction tuning empowers various domain-specific or task-specific LLMs, including natural language generation evaluation (Jiang et al., 2023), math (Yue et al., 2023; Xu et al., 2023a; Azerbayev et al., 2023), code Luo et al. (2023), music Li et al. (2023a); Deng et al. (2023), and medicine Wang et al. (2023a). Instruction tuning not only tailors the models’ task-specific responsiveness but also bolsters their cross-task generalization capabilities, thus enhancing performance across various dynamic application contexts (Wei et al., 2021; Sanh et al., 2022; Mishra et al., 2022; Wang et al., 2022b). Recent studies have broadened the scope of instruction tuning to encompass a wider array of general tasks, notably incorporating input from users of language models (Ouyang et al., 2022; Peng et al., 2023).
@@
Self-Improvement of LLMs In fine-tuning LLMs, the availability of extensive, high-quality instructional data is crucial. Presently, the generation of such data mainly relies on human manual annotation, a labor-intensive method that lacks scalability for future data augmentation. An alternative approach involves deriving instructional data from more advanced LLMs (Taori et al., 2023b), exemplified by extracting authentic instruction-response sets from dialogues within the GPT series models (Wang et al., 2022a). A more refined technique utilizes the Self-Instruction framework, autonomously generating additional instructional data from initial seed data. Combined with the Ada-Instruct or the Evol-Instruct framework, this approach can transform basic instructions into complex ones, specifically tailored for distinct tasks (Cui and Wang, 2023; Luo et al., 2023).
Nevertheless, these instruction generation methodologies all require a robust teacher.The ultimate potential of the model is limited by the teacher’s expertise or resource expenditure (Li et al., 2023c). To overcome this limitation, the SPIN (Chen et al., 2024) framework incorporates a self-play mechanism, It generates training data from previous iterations, refining its strategy by distinguishing between responses generated autonomously and those derived from human-annotated data. This gradual process elevates the LLM from a nascent model to a robust one. Considering the vast amount of knowledge present in web text, Humpback Li et al. (2023c) introduces a technique based on Instruction Backtranslation. This method allows a base model to independently utilize vast amounts of unlabeled data to generate a high-quality instruction tuning dataset. However, empirical findings indicate that the effectiveness of this method is still constrained by the seed model’s performance and its ability to discern truly high-quality data.
@@ 3.3 Refinement of Candidate Labeled Data Using the Primary Chat Model
4 Experiments
4.1 Experimental Setup
4.1.1 Base Model & Finetuning
In our experiments, we utilize the Yi model with 6B parametershttps://huggingface.co/01-ai/Yi-6B, developed by 01.AI, as our foundational language model for fine-tuning. Renowned for its proficiency in both English and Chinese, the Yi series has shown impressive results on global benchmarks like the AlpacaEval Leaderboard (Dubois et al., 2023; Li et al., 2023d) and SuperCLUE (Xu et al., 2023b).
The fine-tuning process is carried out using varying sizes of our high-quality, instructionally curated dataset. This phase is executed on a computing setup with 32 Nvidia A800 GPUs, amounting to a total of 192 GPU hours. We adopt a learning rate of 2e-5 and a batch size of 16, aiming for an optimal balance between computational efficiency and model performance. The generation temperature is set to 0 to ensure precise output generation.
@@ 4.1.2 Baselines
For the Kun datasethttps://huggingface.co/datasets/m-a-p/COIG-Kun, we annotated command data from three sources: Wudao, Wanjuan, and Skypile. Quantitative details of this augmented dataset are provided in Figure 6. In evaluating the performance of Kun, our study contrasts it with data curated from four prominent Chinese open-sourced datasets, including COIG (Zhang et al., 2023; BAAI, 2023a, b), OL-CC (OL-CC, 2023), and BELLE (Ji et al., 2023). These datasets are unique in their composition and focus, providing a comprehensive basis for comparison.
@@ 4.2.1 Statistics and Diversity
Our work involve the purification of approximately 377,592 high-quality instruction-output pairs from the Wudao, Wanjuan, and Skypile datasets. This substantial dataset provide a diverse and rich foundation for our study.
We analyze a 10% subset of instructions from the past 20 years, revealing significant temporal diversity with 56% of instructions from the recent three years (Figure 9). The variation in instruction and output lengths, analyzed using the Yi-6B model, is shown in Figure 2, reflecting content complexity.
To assess instruction diversity, we categorize them into 24 academic disciplines, 16 industry sectors, and 15 text types as per Wikipediahttps://www.wikipedia.org using the Qwen-72B-Chathttps://huggingface.co/Qwen/Qwen-72B-Chat (Bai et al., 2023). Repeated for accuracy, this categorization highlights the data’s range, as shown in Figure 3, where the top 10 categories in each area signify its broad scope.
Figure 6: Distribution of Data Sources
@@ 4.2.2 Quality Evaluation
To ensure objectivity and reliability, three different evaluators evaluate each instruction-output pair. The consistency rates for the evaluation across the three dimensions of the instructions have all exceeded 90%, and the evaluation of the instruction-response also demonstrates a high rate of consistency. This results demonstrate a significant degree of consistency in their judgments. Further details on evaluating identical ratings are presented in Figure 7 and Table 2. Examples that demonstrate the process of assessing can be found in Figure 10
@@
4.3 Result Analysis
As indicated in Table 1, the instruction quality across all sources is consistently high, suggesting effective formulation and clarity in conveying their purposes. However, the output quality varies more noticeably among the sources. While some sources like Wanjuan exhibite a high percentage of "Excellent" outputs, others such as Skypile demonstrate a more diverse distribution of output quality.
This section presents the analysis of our experiment results, encompassing human evaluation outcomes and performance on standard benchmarks.
@@ 4.3.1 Human Evaluation
Figure 8: Heatmap of Model Comparative Win Rates in Human Evaluations
@@ A.2 Data Filtering Rules
  • The length is too short (length <= 4).
@@ A.4 More showcases of human evaluation results
Further analysis is provided in Figures 11, which depict the human preference evaluation win rates of Kun-52k and the mixed model, respectively, when compared with other models. Notably, the mixed model demonstrates a higher win rate than Kun-52k, indicating its enhanced effectiven

### v3

@@ Abstract
In this paper, we introduce Kun [^1] The dataset is named Kun as Chinese pronunciation of Humpback Li et al. (2023c)., a novel approach for creating high-quality instruction-tuning datasets for large language models (LLMs) without relying on manual annotations. Adapting a self-training algorithm based on instruction back-translation and answer polishment, Kun leverages unlabelled data from diverse sources such as Wudao, Wanjuan, and SkyPile to generate a substantial dataset of over a million Chinese instructional data points. This approach presents a novel departure from traditional methods by using a self-curation process to refine and select the most effective instruction-output pairs. Our experiments with the 6B-parameter Yi model across various benchmarks demonstrate Kun’s robustness and scalability. Our method’s core contributions lie in its algorithmic advancement, which enhances data retention and clarity, and its innovative data generation approach that substantially reduces the reliance on costly and time-consuming manual annotations. This methodology presents a scalable and efficient solution for improving the instruction-following capabilities of LLMs, with significant implications for their application across diverse fields.
@@ 1 Introduction
AP is central to Kun’s strategy. It addresses a critical challenge in the Humpback (Li et al., 2023c) method, where raw data, once labeled, are directly used in instruction datasets. The unscreened raw data often mismatches between instructions and responses, as raw data may not inherently align with the instructional context. AP refines this raw data, ensuring a tighter correlation between the instructions and responses through a back-translation process. This leads to a dataset where each instruction-output pair is more coherent and contextually relevant. Unlike methods dependent on LLMs (Peng et al., 2023; Taori et al., 2023b; Zheng et al., 2023), Kun offers an independent and scalable approach to instruction-based training.
We opt for the 6B-parameter Yi model due to its open-source nature and dependable performance [^2] https://github.com/01-ai/Yi. Its efficacy is tested and proven across diverse dataset sizes, including widely recognized benchmarks like C-EVAL (Huang et al., 2023) and CMMLU (Li et al., 2023b). To evaluate the performance of the model, we design a comprehensive human evaluation which contains 500 prompts from ShareGPT-zh, covering various tasks. Responses generated by our model are compared with those from other models, showcasing the superiority of our Kun-52k variant. Further details can be found in 4.2.3. Additionally, we evaluate the quality of our dataset, which includes 1,000 instruction-output pairs each from sources like Wudao (Xue et al., 2022), Wanjuan (He et al., 2023), and SkyPile (Wei et al., 2023). This evaluation, focusing on clarity, feasibility, practicality, and alignment, ensures the high quality of our dataset. The key contributions of our work are:
@@ 2 Related Work
Instruction Tuning
Instruction tuning is widely recognized as a key technique for activating LLMs to adhere to human conversational norms.  (Mishra et al., 2022; Wang et al., 2022b; 2023b). Instruction tuning empowers various domain-specific or task-specific LLMs, including natural language generation evaluation (Jiang et al., 2023), math (Yue et al., 2023; Xu et al., 2023a; Azerbayev et al., 2023), code Luo et al. (2023), music Li et al. (2023a); Deng et al. (2023), and medicine Wang et al. (2023a). Instruction tuning not only tailors the models’ task-specific responsiveness but also bolsters their cross-task generalization capabilities, thus enhancing performance across various dynamic application contexts (Wei et al., 2021; Sanh et al., 2022; Mishra et al., 2022; Wang et al., 2022b). Recent studies have broadened the scope of instruction tuning to encompass a wider array of general tasks, notably incorporating input from users of language models (Ouyang et al., 2022; Peng et al., 2023).
@@
Self-Improvement of LLMs
In fine-tuning LLMs, the availability of extensive, high-quality instructional data is crucial. Presently, the generation of such data mainly relies on human manual annotation, a labor-intensive method that lacks scalability for future data augmentation.
An alternative approach involves deriving instructional data from more advanced LLMs (Taori et al., 2023a), exemplified by extracting authentic instruction-response sets from dialogues within the GPT series models (Wang et al., 2022a). A more refined technique utilizes the Self-Instruction framework, autonomously generating additional instructional data from initial seed data. Combined with the Ada-Instruct or the Evol-Instruct framework, this approach can transform basic instructions into complex ones, specifically tailored for distinct tasks (Cui & Wang, 2023; Luo et al., 2023).
Nevertheless, these instruction generation methodologies all require a robust teacher.The ultimate potential of the model is limited by the teacher’s expertise or resource expenditure (Li et al., 2023c). To overcome this limitation, the SPIN (Chen et al., 2024) framework incorporates a self-play mechanism, It generates training data from previous iterations, refining its strategy by distinguishing between responses generated autonomously and those derived from human-annotated data.
This gradual process elevates the LLM from a nascent model to a robust one. Considering the vast amount of knowledge present in web text, Humpback Li et al. (2023c) introduces a technique based on Instruction Backtranslation. This method allows a base model to independently utilize vast amounts of unlabeled data to generate a high-quality instruction tuning dataset. However, empirical findings indicate that the effectiveness of this method is still constrained by the seed model’s performance and its ability to discern truly high-quality data.
@@ 3.3 Refinement of Candidate Labeled Data Using the Primary Chat Model
4 Experiments
4.1 Experimental Setup
4.1.1 Base Model & Finetuning
In our experiments, we utilize the Yi model with 6B parameters [^3] https://huggingface.co/01-ai/Yi-6B, developed by 01.AI, as our foundational language model for fine-tuning. Renowned for its proficiency in both English and Chinese, the Yi series has shown impressive results on global benchmarks like the AlpacaEval Leaderboard (Dubois et al., 2023; Li et al., 2023d) and SuperCLUE (Xu et al., 2023b).
The fine-tuning process is carried out using varying sizes of our high-quality, instructionally curated dataset. This phase is executed on a computing setup with 32 Nvidia A800 GPUs, amounting to a total of 192 GPU hours. We adopt a learning rate of 2e-5 and a batch size of 16, aiming for an optimal balance between computational efficiency and model performance. All the models have been fine-tuned with the same number of update steps.
@@ 4.1.2 Baselines
For the Kun dataset, we annotated command data from three sources: Wudao, Wanjuan, and Skypile. Quantitative details of this augmented dataset are provided in Figure 7. In evaluating the performance of Kun, our study contrasts it with data curated from four prominent Chinese open-sourced datasets, including COIG (Zhang et al., 2023; BAAI, 2023a; b), OL-CC (OL-CC, 2023), and BELLE (Ji et al., 2023). These datasets are unique in their composition and focus, providing a comprehensive basis for comparison.
@@ 4.2.1 Statistics and Diversity
Our work involve the purification of approximately 377,592 high-quality instruction-output pairs from the Wudao, Wanjuan, and Skypile datasets.We analyze a 10% subset of instructions from the past 20 years, revealing significant temporal diversity with 56% of instructions from the recent three years (Figure 8). The variation in instruction and output lengths, analyzed using the Yi-6B model, is shown in Figure 3, reflecting content complexity.
To assess instruction diversity, we categorize them into 24 academic disciplines, 16 industry sectors, and 15 text types as per Wikipedia [^4] https://www.wikipedia.org using the Qwen-72B-Chat [^5] https://huggingface.co/Qwen/Qwen-72B-Chat (Bai et al., 2023). Repeated for accuracy, this categorization highlights the data’s range, as shown in Figure 2, where the top 10 categories in each area signify its broad scope.
Source Consistency%
Clarity 96.87
Feasibility 97.73
Practicality 97.43
ALL 92.13
Figure 5: Heatmap of Model Comparative Win Rates in Human Evaluations. mix-lite:Kun-26k+COIG-lite.mix-one:Kun-26k+COIG-one
@@ 4.2.2 Quality Evaluation
To ensure objectivity and reliability, three different evaluators evaluate each instruction-output pair. The consistency rates for the evaluation across the three dimensions of the instructions have all exceeded 90%, and the evaluation of the instruction-response are also Consistently. This results demonstrate a significant degree of consistency in their judgments. Further details on evaluating identical ratings are presented in Figure 7 and Table 2. Examples that demonstrate the process of assessing can be found in Figure 11
@@
Figure 7: Distribution of Data Sources
As indicated in Table 1, the instruction quality across all sources is consistently high, suggesting effective formulation and clarity in conveying their purposes. However, the output quality varies more noticeably among the sources. While some sources like Wanjuan exhibite a high percentage of ”Excellent” outputs, others such as Skypile demonstrate a more diverse distribution of output quality. This section presents the analysis of our experiment results, encompassing human evaluation outcomes and performance on standard benchmarks.
@@ A.2 Data Filtering Rules
  • The length is too short (length ¡= 4).
@@ A.4 More showcases of human evaluation results
Further analysis is provided in Figures 9, which depict the human preference evaluation win rates of Kun-52k and the mixed model, respectively, when compared with other models. Notably, the mixed model demonstrates a higher win rate than Kun-52k, indicating its enhanced effectiven.
A.5 Prompt